How many grams of nitrogen gas are present in a 65.0 L gas cylinder at 25oC and 1 atm pressure?


1
Expert's answer
2021-09-30T03:25:02-0400

PV = nRT

1×65.0 = n× 8.314 J/mol·K×298

65= n×2477.572

n= 65/2477.572

n= 0.026

Molar mass of nitrogen = 14.0067

= 14.0067×0.026=0.3675g